You open the window to your apartment and toss a solid metal sphere straight up with a speed of 10 m/s. Your apartment is 13m above the ground (assume this is the starting height of the sphere). What is the maximum height the sphere reaches?
Answer #1


max .heighyt Hmax = 13+(u^2/2g) = 13+(100/19.6) = 18.1m from the ground
Answer #2

let us assume highest height reached by ball above window be h

therefore by velocity - distance equation: h= (v^2-u^2)/2g

where v is the final velocity (in this case ball stops at highest point so v=0m/s)

u is initial velocity=10m/s

g is acceleration due to gravity =9.8 m/s^2

thus h =100/19.6=5.102m

now total height above ground = height of window + h

13 + 5.102 =18.102 m